OpenStudy (anonymous):

the radius of the base of a cylnder is 28 and its hight is 48 what is the surface area of this cylnder? answer in terms of pie

OpenStudy (anonymous):

4256\[\pi\]

OpenStudy (anonymous):

since surface area is the area of the two circles 2((pi)r^2) plus the height which is a rectangle wrapped around the circle, the circumfrence of the circle times the height is that area. then add together

OpenStudy (anonymous):

huh????

OpenStudy (anonymous):

lol sorry the surface area of the cylinder in the area of the two circles plus the cylinder itself. area of one circle = 28^2\[\pi\] = 784pi. multiply by two since there are two circles on a cylinder. the tube of the cylinder can be thought of as a rectangle with one side 48 and the other the circumfrence of the circle. which is 56pi. area of two circles = 1568pi plus the area of the rectangle (48*56pi) = 2688pi. surface area = 1568pi + 2688pi = 4256pi

OpenStudy (anonymous):

oooh ok

OpenStudy (anonymous):

hmm why does the pi always go down to the next line :/

OpenStudy (anonymous):

lol idk k knew what it was though

OpenStudy (anonymous):

* not k i meant but i knew what it was though